# Learning Heuristics for Transit Network Design and Improvement with Deep Reinforcement Learning

### v1

@@ Abstract
Transit agencies world-wide face tightening budgets. To maintain quality of service while cutting costs, efficient transit network design is essential. But planning a network of public transit routes is a challenging optimization problem. The most successful approaches to date use metaheuristic algorithms to search through the space of solutions by applying low-level heuristics that randomly alter routes in a network. The design of these low-level heuristics has a major impact on the quality of the result. In this paper we use deep reinforcement learning with graph neural nets to learn low-level heuristics for an evolutionary algorithm, instead of designing them manually. These learned heuristics improve the algorithm’s results on benchmark synthetic cities with 70 nodes or more, and obtain state-of-the-art results when optimizing operating costs. They also improve upon a simulation of the real transit network in the city of Laval, Canada, by as much as 54% and 18% on two key metrics, and offer cost savings of up to 12% over the city’s existing transit network.
@@ Acknowledgements
We thank the Société de Transport de Laval for the map data and transportation data they provided, and the Agence Métropolitaine de Transport for the origin-destination dataset they provided. This data was essential to the experiments presented in this paper. We also gratefully acknowledge the funding they received from NSERC in support of this work. And we thank Joshua Katz, the room-mate of the first author, for acting as a sounding board for ideas and giving feedback on our experimental designs.
@@ 1 INTRODUCTION
The COVID-19 pandemic resulted in declines in transit ridership in cities worldwide (Liu et al., 2020). This has led to a financial crisis for municipal transit agencies, putting them under pressure to reduce transit operating costs (Kar et al., 2022). But if cost-cutting leads to reduced quality of service, ridership may be harmed further, leading to a downward spiral of service quality and income. Agencies find themselves having to do more with less. Thus, the layout of transit networks is key, as a well-planned layout can cut operating costs and improve service quality, making every dollar count. But updating a transit network is not to be undertaken lightly. For rail transit, changing a network’s layout requires infrastructural changes that may be prohibitively costly. And even for bus transit, network redesigns can be costly and disruptive for users and bus drivers. So it is vital to agencies that they get design right the first time. Good algorithms for the Transit Network Design Problem (NDP) can therefore be very useful to transit agencies.
@@
The NDP is the problem of designing a set of transit routes for a city that satisfy one or more objectives, such as meeting all travel demand and minimizing operating costs. It is an NP-complete problem that has commonalities with the travelling salesman problem (TSP) and vehicle routing problem (VRP), but is more complex than these famously-challenging problems. Furthermore, real-world cities may have hundreds or even thousands of possible stop locations. So analytical optimization approaches are infeasible. For this reason, the most successful algorithms to-date have been metaheuristic algorithms. These work by repeatedly applying one or more low-level heuristics that randomly modify a solution, and guiding this random search towards more promising solutions over many iterations by means of a metaheuristic such as natural selection (as in evolutionary algorithms) or metallic annealing (as in simulated annealing). Many different metaheuristics for guiding the search have been proposed, as have a range of low-level heuristics for transit network design. While these algorithms can find good solutions in many cases, real-world transit networks are still commonly designed by hand (Durán-Micco and Vansteenwegen, 2022).
@@
The low-level heuristics used in these metaheuristic algorithms randomly alter the solutions under consideration, with different heuristics making different kinds of alterations: one heuristic may randomly select a stop on an existing route and remove it from the route, while another may randomly add a stop at one end of a route, and another may select two stops on a route at random and exchange them. What most low-level heuristics have in common is that they are uniformly random: no aspect of the particular scenario affects the heuristic’s likelihood of selecting one action versus another. We here wish to consider whether a machine learning system could act as a more intelligent heuristic, by learning to use information about the city and the current transit network to select the most promising alterations.
@@
In our prior work (Holliday and Dudek, 2023), we trained a neural net to plan a transit network from scratch, and showed that using this transit network as the initial solution improved the solutions found by two metaheuristic algorithms. That policy assembles a transit route by concatenating shortest paths in the city’s street network that share end points, using the neural net to select the next shortest path to be included at each step. By repeating this process for each route that is needed, it assembles a complete transit network.
@@
Our subsequent work (Holliday and Dudek, 2024) built on this by considering whether the graph neural net (GNN) policy can benefit a metaheuristic algorithm by serving as a learned low-level heuristic. To test this, we implemented the evolutionary algorithm of Nikolić and Teodorović (2013). We compared it with a variant in which we replaced one of its low-level heuristics with a heuristic that applies the GNN policy to plan one new route. The GNN policy was first trained in a similar manner as in Holliday and Dudek (2023), before being applied in the evolutionary algorithm. We found that this learned heuristic considerably improves the algorithm’s performance on cities with 70 nodes or more.
@@
In the present work, we expand on these results in several ways. We present new benchmark results using an improved GNN model and an improved evolutionary algorithm. We also perform ablation studies to analyze the importance of different components of our neural-evolutionary algorithm hybrid. As part of these ablations, we consider a novel unlearned low-level heuristic, where shortest paths with common endpoints are joined uniformly at random instead of according to the GNN policy. Interestingly, we find that in the narrow case where we are only concerned with minimizing passenger travel time, this heuristic outperforms both the baseline heuristics of Nikolić and Teodorović (2013) and our GNN heuristic; in most other cases, however, our GNN heuristic performs better.
@@
We compare the performance of our learned and unlearned heuristics with other results on the widely-used Mandl (Mandl, 1980) and Mumford (Mumford, 2013a) benchmark cities. We find that the unlearned heuristic obtains results comparable to the state of the art when minimizing only the passengers’ travel time, while the GNN heuristic outperforms the previous state of the art when minimizing only the cost to operators by as much as 13%.
@@
Finally, we go beyond our earlier work by applying our heuristics to real data from the city of Laval, Canada, a very large real-world problem instance. We show that our learned heuristic can be used to plan transit networks for Laval that, in simulation, exceed the performance of the city’s existing transit system by a wide margin, for three distinct optimization goals. These results show that our heuristics may allow transit agencies to offer better service at less cost.
@@ 2 BACKGROUND AND RELATED WORK
2.1 Neural Nets for Optimization Problems
As documented by  Bengio et al. (2021), there has recently been growing interest in the application of machine learning techniques to combinatorial optimization (CO) problems such as the TSP.
@@ 2.1 Neural Nets for Optimization Problems
Vinyals et al. (2015) proposed a neural net model called a Pointer Network, and trained it via supervised learning to solve TSP instances. Subsequent work, such as that of Dai et al. (2017), Kool et al. (2019), and Sykora et al. (2020), has built on this. These works use similar neural net models, in combination with reinforcement learning algorithms, to train neural nets to construct CO solutions. These have attained impressive performance on the TSP, the VRP, and other CO problems. Choo et al. (2022) present a hybrid algorithm of Monte Carlo Tree Search and Beam Search that draws better sample solutions for the TSP and Capacitated (CVRP) from a neural net policy like that of Kool et al. (2019). Mundhenk et al. (2021) trains an recurrent neural net (RNN) via reinforcement learning (RL) to construct a starting population of solutions to a genetic algorithm, the outputs of which are used to further train the RNN. Fu et al. (2021) train a model on small TSP instances and present an algorithm that applies the model to much larger instances.
@@
These approaches all belong to the family of “construction” methods, which solve a CO problem by starting with an “empty” solution and adding to it in steps, ending once the solution is complete - for example, when all nodes have been visited in the TSP. The solutions from these neural construction methods come close to the quality of those from specialized TSP algorithms such as Concorde (Applegate et al., 2001), while requiring much less run-time to compute (Kool et al., 2019).
@@
By contrast with construction methods, “Improvement” methods start with a complete solution and repeatedly modify it, searching through the solution space for improvements. These are generally more computationally costly than construction methods but can yield better solutions to CO problems. Evolutionary algorithms belong to this category. Some work has considered training neural nets to choose the search moves to be made at each step of an improvement method (Hottung and Tierney, 2019; Chen and Tian, 2019; d O Costa et al., 2020; Wu et al., 2021; Ma et al., 2021), and Kim et al. (2021) train one neural net to construct a set of initial solutions, and another to modify and improve them. This work has shown impressive performance on the TSP and other CO problems. Our work belongs to this family, in that we train a neural net to solve instances of the transit network design problem, and then use it to modify solutions within an improvement method.
[…]
@@
In most CO problems, it is difficult to find globally optimal solutions but easier gauge the quality of a given solution. As  Bengio et al. (2021) note, this makes reinforcement learning, in which machine learning systems are trained to maximize a measure of reward, a natural fit to CO problems. Most of the work cited in this section uses RL methods to train neural net models. We do the same here, applying a method similar to that of Kool et al. (2019) to the NDP.
@@ 2.2 Optimization of Public Transit
The most widely-used metaheuristics for the NDP have been genetic algorithms, simulated annealing, and ant-colony optimization, along with hybrids of these methods (Guihaire and Hao, 2008; Kepaptsoglou and Karlaftis, 2009; Durán-Micco and Vansteenwegen, 2022; Yang and Jiang, 2020; Hüsselmann et al., 2023), achieving impressive results. Recent work has also shown other metaheuristics can be used with success, such as sequence-based selection hyper-heuristics (Ahmed et al., 2019), beam search (Islam et al., 2019), and particle swarms (Lin and Tang, 2022). Many different low-level heuristics have been applied within these metaheuristic algorithms, but most have in common that they select among possible neighbourhood moves uniformly at random.
One exception is Hüsselmann et al. (2023). For two heuristics, the authors design a simple model of how each modification the heuristic could make would affect solution quality. They use this model to give the different modifications different probabilities of being selected. The resulting heuristics obtain state-of-the-art results. However, their simple model ignores both actions’ effect on passenger trips involving transfers, and the impact of the users’ preference over different parts of the cost function. By contrast, the method we propose learns a model of actions’ impact to assign probabilities to them, and does so based on a richer set of inputs.
While neural nets have often been used for predictive problems in urban mobility (Xiong and Schneider, 1992; Rodrigue, 1997; Chien et al., 2002; Jeong and Rilett, 2004; Çodur and Tortum, 2009; Li et al., 2020) and for other transit optimization problems such as scheduling, passenger flow control, and traffic signal control (Zou et al., 2006; Ai et al., 2022; Yan et al., 2023; Jiang et al., 2018; Wang et al., 2024), they have not often been applied to the NDP, and neither has RL. Two recent examples are Darwish et al. (2020) and Yoo et al. (2023). Both use RL to design routes and a schedule for the Mandl benchmark (Mandl, 1980), a single small city with just 15 transit stops, and both obtain good results. Darwish et al. (2020) use a GNN approach inspired by Kool et al. (2019); in our own work we experimented with a nearly identical approach to Darwish et al. (2020), but found it did not scale beyond very small instances. Meanwhile, Yoo et al. (2023) uses tabular RL, a type of approach which is known to scale poorly to large sizes of problem. Both of these approaches also require a new model to be trained on each problem instance. Our approach, by contrast, is able to find good solutions for realistically-sized NDP instances of more than 600 nodes, and can be applied to problem instances unseen during model training.
In our prior work (Holliday and Dudek, 2023), we trained a GNN to construct a transit network one route at a time, and used the GNN to provide initial solutions to two NDP improvement algorithms: the evolutionary algorithm of Nikolić and Teodorović (2013) and the hyperheuristic algorithm of Ahmed et al. (2019). We found that this improved the quality of the solutions found by both methods. In the present work, we consider a similar GNN architecture and training procedure, but we use it as one of the low-level heuristics that is applied at each step of the improvement algorithm.
@@ 3 TRANSIT NETWORK DESIGN PROBLEM
Figure 1: An example city graph with ten numbered nodes and three routes. Street edges are black, routes are in colour, and two example demands are shown by dashed red lines. The edges of the three routes form a subgraph of the street graph $(\mathcal{N},\mathcal{E}_{s})$. All nodes are connected by this subgraph, so the three routes form a valid transit network. The demand between nodes 2 and 5 and between 0 and 6 can be satisfied directly by riding on the blue line, and the demand from 7 to 4 by the orange line. Meanwhile, the demand from 3 to 9 requires one transfer: passengers must ride the orange line from node 3 to 8, and then the green line from node 8 to 9.
@@
This is comprised of a set $\mathcal{N}$ of $n$ nodes, representing candidate stop locations; a set $\mathcal{E}_{s}$ of street edges $(i,j,\tau_{ij})$ connecting the nodes, with weights $\tau_{ij}$ indicating drive times on those streets; and an $n\times n$ Origin-Destination (OD) matrix $D$ giving the travel demand (in number of trips) between every pair of nodes in $\mathcal{N}$. The goal is to propose a transit network, which is a set of routes $\mathcal{R}$ that minimize a cost function $C:\mathcal{C},\mathcal{R}\rightarrow\mathbb{R}^{+}$, where each route $r$ is a sequence of nodes in $\mathcal{N}$. The transit network $\mathcal{R}$ is subject to the following constraints:
@@
    $\mathcal{R}$ must be connected, providing some path over transit between every pair of nodes $(i,j)\in\mathcal{N}$ for which $D_{ij}>0$.
@@
    Every route $r\in\mathcal{R}$ must obey the limit on number of stops per route: $MIN\leq|r|\leq MAX$, where $MIN$ and $MAX$ are parameters set by the user.
@@
    A route $r\in\mathcal{R}$ may not contain cycles; each node $i\in\mathcal{N}$ can appear in $r$ at most once.
We here deal with the symmetric NDP, that is: $D=D^{\top}$, $(i,j,\tau_{ij})\in\mathcal{E}_{s}$ iff. $(j,i,\tau_{ij})\in\mathcal{E}_{s}$, and all routes are traversed both forwards and backwards by vehicles on them. An example city graph with a transit network is shown in Figure 1.

### v2

@@ Abstract
Transit agencies world-wide face tightening budgets. To maintain quality of service while cutting costs, efficient transit network design is essential. But planning a network of public transit routes is a challenging optimization problem. The most successful approaches to date use metaheuristic algorithms to search through the space of possible transit networks by applying low-level heuristics that randomly alter routes in a network. The design of these low-level heuristics has a major impact on the quality of the result. In this paper we use deep reinforcement learning with graph neural nets to learn low-level heuristics for an evolutionary algorithm, instead of designing them manually. These learned heuristics improve the algorithm’s results on benchmark synthetic cities with 70 nodes or more, and obtain state-of-the-art results when optimizing operating costs. They also improve upon a simulation of the real transit network in the city of Laval, Canada, by as much as 54% and 18% on two key metrics, and offer cost savings of up to 12% over the city’s existing transit network.
@@ Acknowledgements
We thank the Société de Transport de Laval for the map data and transportation data they provided, and the Agence Métropolitaine de Transport of Quebec for the origin-destination dataset they provided. This data was essential to the experiments presented in this paper. We also gratefully acknowledge the funding we received from NSERC in support of this work. And we thank Joshua Katz, the room-mate of the first author, for acting as a sounding board for ideas and giving feedback on our experimental designs.
@@ 1 INTRODUCTION
The COVID-19 pandemic resulted in declines in transit ridership in cities worldwide (Liu et al., 2020). This has led to a financial crisis for municipal transit agencies, putting them under pressure to reduce transit operating costs (Kar et al., 2022). But if cost-cutting leads to reduced quality of service, ridership may be harmed further, leading to a downward spiral of service quality and income. Agencies find themselves having to do more with less. Thus, the layout of transit networks is key, as a well-planned layout can cut operating costs and improve service quality. But changes to a transit network are not to be made lightly. For rail transit, it can require infrastructural changes that may be prohibitive. And even for bus transit, network redesigns can be costly and disruptive for users and bus drivers. So it is vital to agencies that they get network design or modification right the first time. Good algorithms for the transit network design problem (NDP) can therefore be very useful to transit agencies.
@@
The NDP is the problem of designing a set of transit routes for a city that satisfy one or more objectives, such as meeting all travel demand and minimizing operating costs. It is an NP-complete problem that has commonalities with the travelling salesman problem (TSP) and vehicle routing problem (VRP), but is more complex than these famously-challenging problems. Since real-world cities typically have hundreds or even thousands of possible stop locations, analytical optimization approaches are infeasible. For this reason, the most successful approaches to the NDP to-date have been metaheuristic algorithms. Most such algorithms work by repeatedly applying one or more low-level heuristics that randomly modify a network, guiding this random walk towards better networks over many iterations by means of a metaheuristic such as natural selection (as in evolutionary algorithms) or metallic annealing (as in simulated annealing). Many different metaheuristics for guiding the search have been proposed, as have a range of low-level heuristics for transit network design specifically. While these algorithms can find good solutions in many cases, real-world transit networks are still commonly designed by hand (Durán-Micco and Vansteenwegen, 2022).
@@
The different low-level heuristics used in these metaheuristic algorithms make different kinds of alterations: one heuristic may randomly select a stop on an existing route and remove it from the route, while another may randomly add a stop at one end of a route, and another may select two stops on a route at random and exchange them. What most low-level heuristics have in common is that they are uniformly random: no aspect of the particular network or city affects the heuristic’s likelihood making one change versus another. We here wish to consider whether a machine learning system could act as a more intelligent heuristic, by learning to use information about the city and the current transit network to make only the most promising changes.
@@
In our prior work (Holliday and Dudek, 2023), we trained a graph neural net (GNN) to plan a transit network from scratch, and showed that using this to generate an initial transit network improved the quality of networks found by two metaheuristic algorithms. That policy assembled transit routes by concatenating shortest paths in the city’s street network that share end-points, using the neural net to select the next shortest path to be included at each step. By repeating this process for each route that is needed, it assembled a complete transit network.
@@
Our subsequent work (Holliday and Dudek, 2024) built on this by considering whether the GNN can benefit a metaheuristic algorithm by serving as a learned low-level heuristic. To test this, we modified an existing evolutionary algorithm (Nikolić and Teodorović, 2013), replacing one of its low-level heuristics with a heuristic that applies the GNN to plan one new route. We found that this learned heuristic considerably improved the algorithm’s performance on cities with 70 nodes or more.
@@
In the present work, we expand on those results in several ways. We present new results obtained with an improved GNN model and an improved evolutionary algorithm. We also perform ablation studies to analyze the importance of different components of our neural-evolutionary hybrid algorithm. As part of these ablations, we consider a novel unlearned low-level heuristic, where shortest paths with common end-points are joined uniformly at random instead of according to the GNN’s output. Interestingly, we find that in the narrow case where we are only concerned with minimizing passenger travel time, this heuristic outperforms both the baseline heuristics of Nikolić and Teodorović (2013) and our GNN heuristic; in most other cases, however, our GNN heuristic performs better.
@@
We compare the performance of our learned and unlearned heuristics with other results on the widely-used Mandl (Mandl, 1980) and Mumford (Mumford, 2013a) benchmark cities. We find that the unlearned heuristic obtains results comparable to the state of the art when minimizing only the passengers’ travel time, while the GNN heuristic outperforms the previous state of the art by as much as 13% when minimizing only the cost to operators.
@@
Finally, we go beyond our earlier work by applying our heuristics to real data from the city of Laval, Canada, a very large real-world problem instance. We show that our learned heuristic can be used to plan transit networks for Laval that, in simulation, exceed the performance of the city’s existing transit system by a wide margin, for three distinct optimization goals. These results show that learned GNN heuristics may allow transit agencies to offer better service at less cost.
@@ 2 BACKGROUND AND RELATED WORK
2.1 Deep Learning for Optimization Problems
Deep learning refers to machine learning techniques that involve “deep” artificial neural nets - that is, neural nets with many hidden layers. Deep reinforcement learning refers to the use of deep neural nets in reinforcement learning (RL). RL is a branch of machine learning in which a learning system, such as a neural net, chooses actions and receives a numerical “reward” for each action, and learns to choose actions that maximize total reward. As documented by  Bengio et al. (2021), there is growing interest in the application of deep learning, and deep RL specifically, to combinatorial optimization (CO) problems such as the TSP.
@@ 2.1 Deep Learning for Optimization Problems
Vinyals et al. (2015) proposed a deep neural net model called a Pointer Network, and trained it via supervised learning to solve TSP instances. Subsequent work, such as that of Dai et al. (2017), Kool et al. (2019), and Sykora et al. (2020), has built on this. These works use similar neural net models, in combination with reinforcement learning algorithms, to train neural nets to construct CO solutions. These have attained impressive performance on the TSP, the VRP, and other CO problems. More recently, Mundhenk et al. (2021) train an recurrent neural net (RNN) via RL to construct a starting population of solutions to a genetic algorithm, the outputs of which are used to further train the RNN. Fu et al. (2021) train a model on small TSP instances and present an algorithm that applies the model to much larger instances. Choo et al. (2022) present a hybrid algorithm of Monte Carlo Tree Search and Beam Search that draws better sample solutions for the TSP and Capacitated (CVRP) from a neural net policy like that of Kool et al. (2019).
@@
These approaches all belong to the family of “construction” methods, which solve a CO problem by starting with an “empty” solution and adding to it until it is complete - for example, in the TSP, this would mean constructing a path one node at a time, starting with an empty path and stopping once the path includes all nodes. The solutions from these neural construction methods come close to the quality of those from specialized algorithms such as Concorde (Applegate et al., 2001), while requiring much less run-time to compute (Kool et al., 2019).
By contrast with construction methods, “improvement” methods start with a complete solution and repeatedly modify it, searching through the solution space for improvements. In the TSP example, this might involve starting with a complete path, and swapping pairs of nodes in the path at each step to see if the path is shortened. Improvement methods are generally more computationally costly than construction methods but can yield better results. Evolutionary algorithms belong to this category.
@@
Some work has considered training neural nets to choose the search moves to be made at each step of an improvement method (Hottung and Tierney, 2019; Chen and Tian, 2019; d O Costa et al., 2020; Wu et al., 2021; Ma et al., 2021), and Kim et al. (2021) train one neural net to construct a set of initial solutions, and another to modify and improve them. This work has shown impressive performance on the TSP and other CO problems. Our work belongs to this family, in that we train a neural net to solve instances of the transit network design problem, and then use it to modify solutions within an improvement method.
In most of the above work, the neural net models used are graph neural nets, a type of neural net model that is designed to operate on graph-structured data (Bruna et al., 2013; Kipf and Welling, 2016; Defferrard et al., 2016; Duvenaud et al., 2015). These have been applied in many other domains, including the analysis of large web graphs (Ying et al., 2018), the design of printed circuit boards (Mirhoseini et al., 2021), and the prediction of chemical properties of molecules (Duvenaud et al., 2015; Gilmer et al., 2017). An overview of GNNs is provided by Battaglia et al. (2018). Like many CO problems, the transit network design problem lends itself to being described as a graph problem, so we use GNN models here as well.
@@
In most CO problems, it is difficult to find globally optimal solutions but easier gauge the quality of a given solution. As  Bengio et al. (2021) note, this makes reinforcement learning a natural fit to CO problems. Most of the work cited in this section uses RL methods to train neural net models. We do the same here, applying a method similar to that of Kool et al. (2019) to the NDP.
@@ 2.2 Optimization of Public Transit
The most widely-used metaheuristics for the NDP have been genetic algorithms, simulated annealing, and ant-colony optimization, along with hybrids of these methods (Guihaire and Hao, 2008; Kepaptsoglou and Karlaftis, 2009; Durán-Micco and Vansteenwegen, 2022; Yang and Jiang, 2020; Hüsselmann et al., 2023). Recent work has also shown other metaheuristics can be used with success, such as sequence-based selection hyper-heuristics (Ahmed et al., 2019), beam search (Islam et al., 2019), and particle swarms (Lin and Tang, 2022). Many different low-level heuristics have been applied within these metaheuristic algorithms, but most have in common that they select among possible neighbourhood moves uniformly at random.
One exception is Hüsselmann et al. (2023). For two heuristics, the authors design a simple model of how each change the heuristic could make would affect solution quality. They use this model to give the different changes different probabilities of being selected. The resulting heuristics obtain state-of-the-art results. However, their simple model ignores passenger trips involving transfers, and the impact of the user’s preferences over different parts of the cost function. By contrast, the method we propose learns a model of changes’ impacts to assign probabilities to those changes, and does so based on a richer set of inputs.
While neural nets have often been used for predictive problems in urban mobility (Xiong and Schneider, 1992; Rodrigue, 1997; Chien et al., 2002; Jeong and Rilett, 2004; Çodur and Tortum, 2009; Li et al., 2020) and for other transit optimization problems such as scheduling, passenger flow control, and traffic signal control (Zou et al., 2006; Ai et al., 2022; Yan et al., 2023; Jiang et al., 2018; Wang et al., 2024), they have not often been applied to the NDP, and neither has RL. Two recent examples are Darwish et al. (2020) and Yoo et al. (2023). Both use RL to design routes and a schedule for the Mandl benchmark (Mandl, 1980), a single small city with just 15 transit stops, and both obtain good results. Darwish et al. (2020) use a GNN approach inspired by Kool et al. (2019); in our own work we experimented with a nearly identical approach to Darwish et al. (2020), but found it did not scale beyond very small instances. Meanwhile, Yoo et al. (2023) use tabular RL, a type of approach which is practical only for small problem sizes. Both of these approaches also require a new model to be trained on each problem instance. Our approach, by contrast, is able to find good solutions for NDP instances of more than 600 nodes, and can be applied to instances unseen during model training.
@@ 3 TRANSIT NETWORK DESIGN PROBLEM
Figure 1: An example city graph with ten numbered nodes and three routes. Street edges are black, routes are in colour, and demands are shown by dashed red lines. The edges of the three routes form a subgraph of the street graph $(\mathcal{N},\mathcal{E}_{s})$. All node-pairs with demand are connected by this subgraph, so the three routes form a valid transit network. The demand between nodes 2 and 5 and between 0 and 6 can be satisfied directly by riding on the blue line, and the demand from 7 to 4 by the orange line. But the demand from 3 to 9 requires one transfer: passengers must ride the orange line from node 3 to 8, and then the green line from node 8 to 9.
@@
This is comprised of a set $\mathcal{N}$ of $n$ nodes, representing candidate stop locations; a set $\mathcal{E}_{s}$ of street edges $(i,j,\tau_{ij})$ connecting the nodes, with weights $\tau_{ij}$ indicating drive times on those streets; and an $n\times n$ Origin-Destination (OD) matrix $D$ giving the travel demand, in number of trips, between every pair of nodes in $\mathcal{N}$. The goal is to propose a transit network, which is a set of routes $\mathcal{R}$ that minimizes a cost function $C:\mathcal{C},\mathcal{R}\rightarrow\mathbb{R}^{+}$, where each route $r\in\mathcal{R}$ is a sequence of nodes in $\mathcal{N}$. The transit network $\mathcal{R}$ is subject to the following constraints:
@@
    $\mathcal{R}$ must satsify all demand, providing some path over transit between every pair of nodes $(i,j)\in\mathcal{N}$ for which $D_{ij}>0$.
@@
    All routes $r\in\mathcal{R}$ must obey the same limits on its number of stops: $MIN\leq|r|\leq MAX$, where $MIN$ and $MAX$ are parameters set by the user.
@@
    Routes $r\in\mathcal{R}$ must not contain cycles; each node $i\in\mathcal{N}$ can appear in $r$ at most once.
We here deal with the symmetric NDP, meaning that all demand, streets, and routes are the same in both directions. This means $D=D^{\top}$, $(i,j,\tau_{ij})\in\mathcal{E}_{s}$ iff. $(j,i,\tau_{ij})\in\mathcal{E}_{s}$, and all routes are traversed both forwards and backwards by vehicles on them. An example city graph with a transit network is shown in Figure 1.